# II. Carbon-enhanced metal-poor starsthanks: All M3DIS 3D RHD CEMP models are available under https://nlte.mpia.de

### v1

@@ 3.2 Spectrum Synthesis of the CH G-band
We synthesize the CH G-band for all models listed in Tab. 2, including their scaled-solar counterpart, and for comparison, the respective 1D MARCS models with the same stellar parameters.
@@
In Fig. 4 we show two representative synthetic G-band profiles for two model atmospheres, MS (top) and sub-giant (bottom), at $\rm[Fe/H]=-5$ and $\rm[C/Fe]=+3$. Independent of stellar parameters, it is evident that the 3D CEMP model (black, solid lines) leads to significantly stronger CH features than the 1D model (red, solid lines) at the same $\rm A(C)$. The 3D scaled-solar model (blue, dashed lines) on the other hand produces an even stronger G-band as the 3D CEMP model (for the same Teff, $\rm\log(g)$, and metallicity) and hence shows the largest departure from 1D HE. This is reflected in the equivalent width ($\rm EW$).
@@
Figure 5: CH and C II number densities as a function of optical depth. Solid, blue lines correspond to the 3D scaled-solar model, blue crosses to the 3D CEMP model. The 1D MARCS model is shown with red, solid lines. For all models, CH number densities are shown as opaque lines, while C II number densities are shown as thick lines with higher transparency.
@@
Figure 6: Abundance corrections to A(C) derived from the equivalent width of the CH G-band between $\rm 4\,297\ and\ 4\,303\AA$ . Left (panel A): Final abundance correction $\rm\Delta A(C)=A(C)_{3D\ CEMP}-A(C)_{1D}$ from $\rm 1D\ HE\ MARCS$ to $\rm 3D\ CEMP$ models. Centre (panel B): $\rm A(C)_{3D\ CEMP}-A(C)_{3D\ \text{scaled-solar}}$. Right (panel C): $\rm A(C)_{3D\ \text{scaled-solar}}-A(C)_{1D}$.
@@
These strong differences in the G-band strength are due to the differences in the atmospheric structure. To explore this connection in more detail, we show the average CH number density, $\rm\log n_{CH}$, in the atmosphere of the sub-giant model at $\rm[Fe/H]=-5$ as a function of optical depth in Fig. 5. We additionally include the density of singly-ionized carbon (C II). The CH number density can best be understood by linking their evolution with optical depth to the differences in the atmospheric structures as presented in Fig. 1, panel D. Below $\rm\log\tau_{500}=-1$ the atmosphere is dropping sharply in temperature – in 3D as well as in 1D – which favours the formation of CH molecules at the expense of C II, which is recombining with free electrons in the successively cooler and less dense environment. Towards the upper boundary the temperature in the 1D HE model stays roughly constant, while the density is dropping exponentially. In chemical equilibrium the CH number density is proportional to $\rm n_{C}$ – as well as $\rm n_{H}$ (Eq. 1) – and hence $\rm n_{CH}$ drops proportional to $\rm\rho^{2}$. This explains why the CH number density decreases significantly in the 1D model, despite the temperature staying constant. The C II number density on the other hand is only linearly dependent on density, and additionally is affected by the balance of $\rm H^{-}$ and C II dissociation energies, which is why the trend is less obvious.
@@
While this proportionality is also true for the 3D model, the behaviour in the atmosphere is entirely different. This is because the 3D model is not in radiative equilibrium – it is dominated by adiabatic cooling – which means the temperature drops as a function of decreasing optical depth. In an increasingly colder environment the conditions for CH formation improve drastically, which is why $\rm n_{CH}$ increases in both 3D models and departures between 3D and 1D become significant. Comparing 3D CEMP (blue crosses) and 3D scaled-solar (blue, solid lines) it may be noted that there are significantly more CH molecules present in the upper layers of the scaled-solar model, while there is only very little C II, compared to the CEMP model. Looking again at Fig. 1 this difference can be attributed to the additional heating induced by the C-enhancement in the CEMP atmosphere, which causes it to be hotter and hence less favourable for CH formation.
@@
3D RHD models are much, up to 1000-1000 K, colder in atmospheric layers above $\log\tau_{500}\lesssim-2.5$. Because of the strong C-enhancement, the cores of strongest lines in the G-band form in regions around $\rm\log\tau_{500}=-3.5$, as indicated by the horizontal lines in Fig. 2 and the $\rm n_{CH}$ extrema in Fig. 5. Due to much lower temperatures, the formation of molecules is more efficient, and due to larger molecular number densities the absorption profiles are stronger (Uitenbroek & Criscuoli 2011). The final consequence is that the presence of strong G-band absorption in an observed spectrum is not necessarily linked to a significant C-enhancement, but can, in part, be associated with strong convection as expected at low metallicities (Collet et al. 2011; Gallagher et al. 2017).
@@
The larger differences seen in the synthetic spectra for the sub-giant (Fig. 4, bottom panel), as compared to the main-sequence star, directly reflect the atmospheric structure distinctions discussed in Sec. 3.1.3. Specifically, while the 1D sub-giant model is colder (enhancing its G-band), the interplay of the non-linear dependence of CH formation on temperature and density, combined with the more pronounced C-enhancement induced re-heating effect in the 3D sub-giant atmosphere, leads to the observed spectral differences relative to 1D.
@@ 3.3 CEMP abundances in 3D
In order to estimate the influence of carbon-enhanced CEMP 3D RHD models on synthetic observables, we perform the analysis as presented in the previous sections for stellar model atmospheres of different metallicities, from $\rm[Fe/H]=-2$ up to $\rm[Fe/H]=-6$. We compare the total EWs in the wavelength range of the G-band obtained in 3D with the corresponding 1D MARCS results, computed using a micro-turbulence of 1 km//s. By varying A(C) in the spectrum synthesis of the 3D RHD models we sample the full curve-of-growth and derive abundance corrections by matching the EWs within the integration limit of 4297 to 4303 Å. We emphasize that this EW-integration of the G-band is not used to obtain abundances of observed spectra directly, but only to quantify the corresponding overall effect on carbon abundances by comparing synthetic spectra. Other procedures, such as direct fitting of lines (Popa et al. 2023) can be used instead.
@@
As explained in the previous section, $\rm 3D-1D$ abundance corrections are larger (in modulus, or amplitude) for the sub-giant star. At $\rm[Fe/H]=-6$ they are large and negative, and reach up to $\rm-0.7\ dex$ with C-enhancement in the atmosphere, and $\rm-0.9\ dex$ without. Corrections decrease with increasing metallicity and are smaller, $\rm\Delta A(C)\approx-0.2\ dex$, at $\rm[Fe/H]=-2$. The $\rm 3D\ CEMP-1D$ corrections for the hotter MS star range from $\rm\sim-0.6\ dex$ at [Fe/H] =6=-6 to $\rm\sim-0.2\ dex$ at [Fe/H] =2=-2. In contrast, the effect of carbon enhancement in 3D models is positive (panel B). At [Fe/H] =2=-2, 3D scaled-solar and 3D CEMP G-band strengths in the sub-giant models are almost identical, which is why $\rm A(C)_{3D\ CEMP}-A(C)_{3D\ \text{scaled-solar}}\lesssim 0.05\ dex$. The differences increase with decreasing metallicity and decreasing surface gravity, so that $\rm A(C)_{3D\ CEMP}-A(C)_{3D\ \text{scaled-solar}}=\rm+0.2\ dex$ for the sub-giant model with $\rm[Fe/H]=-6$. As discussed in Sec. 3.1.2, the effect of C-enhancement on the hotter MS model is overall smaller. The corrections are on the order of $\rm\lesssim+0.1\ dex$, and show only a small deviation with metallicity. Compared to the significant 3D effect, the C-enhancement in hot, low metallicity MS models is hence negligible.
@@
Figure 7: [C/Fe] against [Fe/H] for 3D CEMP and 1D measurements of Galactic stars from a MS + sub-giant sub-sample of the SAGA database (see text). The red, dashed line represents the mean, uncorrected 1D and the blue, solid line the 3D CEMP-corrected data. Shaded regions correspond to standard deviations of the observed sample in the respective metallicity bin $\rm\pm 0.5\ dex$. The grey dotted line represents the dividing CEMP classification abundance of $\rm[C/Fe]=0.7$.
@@
Figure 8: Yoon-Beers (Yoon et al. 2016) diagram, A(C) against [Fe/H] for the 3D-corrected measurements for Galactic stars from the literature (see text). The grey dotted line represents the dividing CEMP classification abundance of $\rm[C/Fe]=0.7$. Coloured lines show the major axis of the ellipse representing the distribution of different CEMP groups. In 1D (dashed lines), ellipses are drawn similar to Yoon et al. (2016), their Fig. 1. In 3D (solid lines) ellipses are obtained from fitting a Gaussian mixture model to the data.
@@
It is interesting to analyse the impact of 3D CEMP models in the context of the observed CEMP distribution of Galactic stars. We therefore apply our corrections to a sample of MS and turn-off stars compiled from the SAGA database (Suda et al. 2008, 2011, 2017) to explore whether 3D models have an impact on statistics of CEMP stars in different Galactic metallicity regimes. We apply $\rm A(C)_{3D\ CEMP}-A(C)_{1D}$ carbon abundance corrections to the observed data, based on the closest 3D model in terms of $\rm T_{eff}$ and $\rm log(g)$, and we interpolate linearly in metallicity. We then split the stellar sample into bins of $\rm\pm 0.5\ dex$ in metallicity and compute mean and standard deviation ($\sigma$) in each of them. The result is shown in Fig. 7. As expected, the differences between the 3D-corrected and 1D distributions increase with decreasing [Fe/H]. At $\rm[Fe/H]=-4$ the mean $\rm[C/Fe]$ of CEMP stars is $\rm\approx 1.1$, $\rm 0.5\ dex$ lower than the $\rm[C/Fe]\approx 1.6$ that the 1D data suggest. Towards $\rm[Fe/H]=-2$ the differences decrease to $\rm\approx 0.2\ dex$. We note that at metallicities above 5-5 the lower limit of the 1-$\sigma$ interval already lies below the CEMP defining line of $\rm[C/Fe]=0.7$ (indicated by the grey, dotted line) in 3D, which implies that there is a significant impact on the overall abundance distribution in this metallicity regime. Below $\rm[Fe/H]=-5$ the corrections are larger, however because of the overall higher $\rm[C/Fe]$ abundances found in these stars the impact on the distribution of CEMP stars is expected to be small.
@@
To investigate Galactic CEMP populations further, we present the same data in the so-called “Yoon-Beers” diagram (Yoon et al. 2016), which is shown in Fig. 8. In this figure, we indicate the location of CEMP sub-groups I, II, and III – as defined by Yoon et al. (2016), see their Fig. 1 – as coloured lines representing the major axis of the corresponding bivariate normal distribution ellipses. In 1D, the ellipses are adopted from Yoon et al. (2016), while in 3D we fit a Gaussian mixture model to our data. In addition to the first identification of CEMP sub-groups, Yoon et al. furthermore showed that group I is mainly populated by CEMP-r/s stars, while group II and III contain mainly CEMP-no stars. Yoon et al. (2016) – and subsequently also Yoon et al. (2018); Chiaki et al. (2017) – furthermore identify different CEMP groups with different formation scenarios, indicating multiple progenitors for CEMP-no stars.
@@
At lower metallicities, differences manifest as a systematic offset of $\rm\sim 0.6\ dex$, shifting down the $\rm A(C)$ plateau (indicated by the horizontal orientation of the red line in the 1D case) significantly. We note that, as a consequence of this plateau, the carbon enhancement in the atmosphere, relative to iron, becomes more significant when considering even lower metallicities. We do not expect that – at least in CEMP stars – G-bands from 3D and 1D models diverge even further at more extreme metallicities, mainly because the carbon enhancement itself will compensate for the loss of metals in exactly the right part of the atmosphere, as demonstrated in the previous sections. This compensation is noticeable as a decreasing metallicity gradient in panel A (compared to panel C) of Fig. 6 already between $\rm[Fe/H]=-4\ and-5$. One may be tempted to use scaled-solar, low metallicity, non-carbon-enhanced models for the analysis of these CEMP stars due to the time investment related to creating 3D models of stellar atmospheres. However, this would lead to an overestimation of the corrections and introduce an additional differential effect with metallicity that causes the A(C) plateau to be artificially tilted. When properly including C-enhancement in the atmosphere computation, this tilt is weaker, but still noticeable mostly due to differences in stellar parameters of individual targets. In Fig. 8 we find that an ellipse tilted by $\rm 10^{\circ}$ with respect to the 1D case best represents our group III sample in 3D.
@@
Because $\rm 3D-1D$ abundance corrections are negative, we find a significant number of stars in the SAGA sample that are – following the classical definition of Aoki et al. (2006) – no longer CEMP stars, as their $\rm[C/Fe]$ drops below $\rm 0.7$. This is most noticeable in CEMP group II, as these stars are located closest to the carbon-normal division in terms of $\rm A(C)$, while still experiencing a large abundance correction due to their relatively low metallicities. Around the onset of the carbon-plateau at $\rm[Fe/H]=-4$ we find that stars at higher $\rm[C/Fe]$ are systematically redistributed towards the abundance threshold, such that close to – and below – the dividing line we find significantly more stars in 3D than in 1D. Overall, there are significantly fewer stars with extreme carbon abundances above $\rm[C/Fe]=1.5$ (see Fig. 7, specifically for $\rm[Fe/H]\geq-4$). The consequence is that that there is a migration of stars at the edges of group I and III towards group II, or even outside the CEMP category. In 3D, we find that group II is best represented by an ellipse that is shifted down by $\rm 0.4\ dex$ with respect to the 1D case.
@@
Figure 9: Cumulative fraction of CEMP-no (solid lines with squares) and CEMP-r/s (dashed lines with circles) stars in 1D (red, left panel) as well as 3D (blue, right panel). The dotted line with crosses shows the total cumulative fraction of CEMP stars in 1D (left) and 3D (right).

### v2

@@ 3.2 Spectrum synthesis of the CH G band
We synthesised the CH G band for all models listed in Table 2, including their scaled-solar counterpart, and for comparison, the respective 1D MARCS models with the same stellar parameters.
@@
In Fig. 4 we show two representative synthetic G-band profiles for two model atmospheres, MS (top) and sub-giant (bottom), at $\rm[Fe/H]=-5$ and $\rm[C/Fe]=+3$. Independent of stellar parameters, it is evident that the 3D CEMP model (solid black lines) leads to significantly stronger CH features than the 1D model (solid red lines) at the same $\rm A(C)$. The 3D scaled-solar model (dashed blue lines) on the other hand produces an even stronger G band as the 3D CEMP model (for the same Teff, $\rm\log(g)$, and metallicity) and hence shows the largest departure from 1D HE. This is reflected in the equivalent width (EW).
@@
Figure 5: CH and C II number densities as a function of optical depth. Solid blue lines correspond to the 3D scaled-solar model, blue crosses to the 3D CEMP model. The 1D MARCS model is shown with solid red lines. For all models, CH number densities are shown as opaque lines, while C II number densities are shown as thick, partially transparent lines.
@@
Figure 6: Abundance corrections to A(C) derived from the EW of the CH G band between $\rm 4\,297\ and\ 4\,303\AA$ . Panel A: Final abundance correction $\rm\Delta A(C)=A(C)_{3D\ CEMP}-A(C)_{1D}$ from $\rm 1D\ HE\ MARCS$ to $\rm 3D\ CEMP$ models. Panel B: $\rm A(C)_{3D\ CEMP}-A(C)_{3D\ \text{scaled-solar}}$. Panel C: $\rm A(C)_{3D\ \text{scaled-solar}}-A(C)_{1D}$.
@@
These strong differences in the G-band strength are due to the differences in the atmospheric structure. To explore this connection in more detail, we show the average CH number density, $\rm\log n_{CH}$, in the atmosphere of the sub-giant model at $\rm[Fe/H]=-5$ as a function of optical depth in Fig. 5. We additionally include the density of singly ionised carbon (C II). The CH number density can best be understood by linking their evolution with optical depth to the differences in the atmospheric structures as presented in Fig. 1, panel D. Below $\rm\log\tau_{500}=-1$ the atmosphere is dropping sharply in temperature – in 3D as well as in 1D – which favours the formation of CH molecules at the expense of C II, which is recombining with free electrons in the successively cooler and less dense environment. Towards the upper boundary the temperature in the 1D HE model stays roughly constant, while the density is dropping exponentially. In chemical equilibrium the CH number density is proportional to $\rm n_{C}$ – as well as $\rm n_{H}$ (Eq. 1) – and hence $\rm n_{CH}$ drops proportional to $\rm\rho^{2}$. This explains why the CH number density decreases significantly in the 1D model, despite the temperature staying constant. The C II number density on the other hand is only linearly dependent on density, and additionally is affected by the balance of $\rm H^{-}$ and C II dissociation energies, which is why the trend is less obvious.
@@
While this proportionality is also true for the 3D model, the behaviour in the atmosphere is entirely different. This is because the 3D model is not in radiative equilibrium – it is dominated by adiabatic cooling – which means the temperature drops as a function of decreasing optical depth. In an increasingly colder environment the conditions for CH formation improve drastically, which is why $\rm n_{CH}$ increases in both 3D models and departures between 3D and 1D become significant. Comparing 3D CEMP (blue crosses) and 3D scaled-solar (solid blue lines), we see that there are significantly more CH molecules present in the upper layers of the scaled-solar model, while there is only very little C II, compared to the CEMP model. Looking again at Fig. 1 this difference can be attributed to the additional heating induced by the C-enhancement in the CEMP atmosphere, which causes it to be hotter and hence less favourable for CH formation.
@@
3D RHD models are much, up to 1000-1000 K, colder in atmospheric layers above $\log\tau_{500}\lesssim-2.5$. Because of the strong C-enhancement, the cores of strongest lines in the G band form in regions around $\rm\log\tau_{500}=-3.5$, as indicated by the horizontal lines in Fig. 2 and the $\rm n_{CH}$ extrema in Fig. 5. Due to much lower temperatures, the formation of molecules is more efficient, and due to larger molecular number densities the absorption profiles are stronger (Uitenbroek & Criscuoli 2011). The final consequence is that the presence of strong G-band absorption in an observed spectrum is not necessarily linked to a significant C-enhancement but can, in part, be associated with strong convection as expected at low metallicities (Collet et al. 2011; Gallagher et al. 2017).
@@
The larger differences seen in the synthetic spectra for the sub-giant (Fig. 4, bottom panel), as compared to the MS star, directly reflect the atmospheric structure distinctions discussed in Sect. 3.1.3. Specifically, while the 1D sub-giant model is colder (enhancing its G band), the interplay of the non-linear dependence of CH formation on temperature and density, combined with the more pronounced C-enhancement induced re-heating effect in the 3D sub-giant atmosphere, leads to the observed spectral differences relative to 1D.
@@ 3.3 CEMP abundances in 3D
In order to estimate the influence of carbon-enhanced CEMP 3D RHD models on synthetic observables, we performed the analysis as presented in the previous sections for stellar model atmospheres of different metallicities, from $\rm[Fe/H]=-2$ up to $\rm[Fe/H]=-6$. We compared the total EWs in the wavelength range of the G band obtained in 3D with the corresponding 1D MARCS results, computed using a micro-turbulence of 1 km//s. By varying A(C) in the spectrum synthesis of the 3D RHD models, we sampled the full curve of growth and derived abundance corrections by matching the EWs within the integration limit of 4297 to 4303 Å. We emphasise that this EW-integration of the G band is not used to obtain abundances of observed spectra directly, but only to quantify the corresponding overall effect on carbon abundances by comparing synthetic spectra. Other procedures, such as a direct fitting of lines (Popa et al. 2023), can be used instead.
@@
As explained in the previous section, $\rm 3D-1D$ abundance corrections are larger (in modulus, or amplitude) for the sub-giant star. At $\rm[Fe/H]=-6$ they are large and negative, and reach up to $\rm-0.7\ dex$ with C-enhancement in the atmosphere, and $\rm-0.9\ dex$ without. Corrections decrease with increasing metallicity and are smaller, $\rm\Delta A(C)\approx-0.2\ dex$, at $\rm[Fe/H]=-2$. The $\rm 3D\ CEMP-1D$ corrections for the hotter MS star range from $\rm\sim-0.6\ dex$ at [Fe/H] =6=-6 to $\rm\sim-0.2\ dex$ at [Fe/H] =2=-2. In contrast, the effect of carbon enhancement in 3D models is positive (panel B). At [Fe/H] =2=-2, 3D scaled-solar and 3D CEMP G-band strengths in the sub-giant models are almost identical, which is why $\rm A(C)_{3D\ CEMP}-A(C)_{3D\ \text{scaled-solar}}\lesssim 0.05\ dex$. The differences increase with decreasing metallicity and decreasing surface gravity, so that $\rm A(C)_{3D\ CEMP}-A(C)_{3D\ \text{scaled-solar}}=\rm+0.2\ dex$ for the sub-giant model with $\rm[Fe/H]=-6$. As discussed in Sect. 3.1.2, the effect of C-enhancement on the hotter MS model is overall smaller. The corrections are of the order of $\rm\lesssim+0.1\ dex$, and show only a small deviation with metallicity. Compared to the significant 3D effect, the C-enhancement in hot, low metallicity MS models is hence negligible.
@@
Figure 7: [C/Fe] against [Fe/H] for 3D CEMP and 1D measurements of Galactic stars from a MS + sub-giant subsample of the SAGA database. The dashed red line represents the mean, uncorrected 1D and the solid blue line the 3D CEMP-corrected data. Shaded regions correspond to standard deviations of the observed sample in the respective metallicity bin $\rm\pm 0.5\ dex$. The dotted grey line represents the dividing CEMP classification abundance of $\rm[C/Fe]=0.7$.
@@
Figure 8: Yoon-Beers (Yoon et al. 2016) diagram of A(C) against [Fe/H] for the 3D-corrected measurements for Galactic stars from the literature. The dotted grey line represents the dividing CEMP classification abundance of $\rm[C/Fe]=0.7$. Coloured lines show the major axis of the ellipse representing the distribution of different CEMP groups. In 1D (dashed lines), ellipses are drawn similar to Yoon et al. (2016), their Fig. 1. In 3D (solid lines) ellipses are obtained from fitting a Gaussian mixture model to the data.
@@
It is interesting to analyse the impact of 3D CEMP models in the context of the observed CEMP distribution of Galactic stars. We therefore applied our corrections to a sample of MS and turnoff stars compiled from the SAGA database (Suda et al. 2008, 2011, 2017) to explore whether 3D models have an impact on statistics of CEMP stars in different Galactic metallicity regimes. We applied $\rm A(C)_{3D\ CEMP}-A(C)_{1D}$ carbon abundance corrections to the observed data, based on the closest 3D model in terms of $\rm T_{eff}$ and $\rm log(g)$, and we interpolated linearly in metallicity. We then split the stellar sample into bins of $\rm\pm 0.5\ dex$ in metallicity and computed the mean and standard deviation ($\sigma$) in each of them. The result is shown in Fig. 7. As expected, the differences between the 3D-corrected and 1D distributions increase with decreasing [Fe/H]. At $\rm[Fe/H]=-4$ the mean $\rm[C/Fe]$ of CEMP stars is $\rm\approx 1.1$, $\rm 0.5\ dex$ lower than the $\rm[C/Fe]\approx 1.6$ that the 1D data suggest. Towards $\rm[Fe/H]=-2$ the differences decrease to $\rm\approx 0.2\ dex$. We note that at metallicities above 5-5 the lower limit of the 1-$\sigma$ interval already lies below the CEMP defining line of $\rm[C/Fe]=0.7$ (indicated by the dotted grey line) in 3D, which implies that there is a significant impact on the overall abundance distribution in this metallicity regime. Below $\rm[Fe/H]=-5$ the corrections are larger; however, because of the overall higher $\rm[C/Fe]$ abundances found in these stars, the impact on the distribution of CEMP stars is expected to be small.
@@
To investigate Galactic CEMP populations further, we present the same data in the so-called Yoon-Beers diagram (Yoon et al. 2016), which is shown in Fig. 8. In this figure, we show the location of CEMP subgroups I, II, and III as defined by Yoon et al. (2016); see their Fig. 1. In 1D, the ellipses are adopted from Yoon et al. (2016), while in 3D we fitted a Gaussian mixture model to our data. In addition to the first identification of CEMP subgroups, Yoon et al. furthermore showed that group I is mainly populated by CEMP-r/s stars, while groups II and III contain mainly CEMP-no stars. Yoon et al. (2016) – and subsequently also Yoon et al. (2018); Chiaki et al. (2017) – furthermore identify different CEMP groups with different formation scenarios, indicating multiple progenitors for CEMP-no stars.
@@
At lower metallicities, differences manifest as a systematic offset of $\rm\sim 0.6\ dex$, shifting down the $\rm A(C)$ plateau (indicated by the horizontal orientation of the red line in the 1D case) significantly. We note that, as a consequence of this plateau, the carbon enhancement in the atmosphere, relative to iron, becomes more significant when considering even lower metallicities. We do not expect that – at least in CEMP stars – G bands from 3D and 1D models diverge even further at more extreme metallicities, mainly because the carbon enhancement itself will compensate for the loss of metals in exactly the right part of the atmosphere, as demonstrated in the previous sections. This compensation is noticeable as a decreasing metallicity gradient in panel A (compared to panel C) of Fig. 6 already between $\rm[Fe/H]=-4\ and-5$. One may be tempted to use scaled-solar, low metallicity, non-carbon-enhanced models for the analysis of these CEMP stars due to the time investment related to creating 3D models of stellar atmospheres. However, this would lead to an overestimation of the corrections and introduce an additional differential effect with metallicity that causes the A(C) plateau to be artificially tilted. When properly including C-enhancement in the atmosphere computation, this tilt is weaker, but still noticeable mostly due to differences in stellar parameters of individual targets. In Fig. 8 we find that an ellipse tilted by $\rm 10^{\circ}$ with respect to the 1D case best represents our group III sample in 3D.
@@
Because $\rm 3D-1D$ abundance corrections are negative, we find a significant number of stars in the SAGA sample that are – following the classical definition of Aoki et al. (2006) – no longer CEMP stars, as their $\rm[C/Fe]$ drops below $\rm 0.7$. This is most noticeable in CEMP group II, as these stars are located closest to the carbon-normal division in terms of $\rm A(C)$, while still experiencing a large abundance correction due to their relatively low metallicities. Around the onset of the carbon-plateau at $\rm[Fe/H]=-4$ we find that stars at higher $\rm[C/Fe]$ are systematically redistributed towards the abundance threshold, such that close to – and below – the dividing line we find significantly more stars in 3D than in 1D. Overall, there are significantly fewer stars with extreme carbon abundances above $\rm[C/Fe]=1.5$ (see Fig. 7, specifically for $\rm[Fe/H]\geq-4$). The consequence is that that there is a migration of stars at the edges of groups I and III towards group II, or even outside the CEMP category. In 3D, we find that group II is best represented by an ellipse that is shifted down by $\rm 0.4\ dex$ with respect to the 1D case.
@@
Figure 9: Cumulative fraction of CEMP-no (solid lines with squares) and CEMP-r/s (dashed lines with circles) stars in 1D (red, left panel) and 3D (blue, right panel). The dotted line with crosses shows the total cumulative fraction of CEMP stars in 1D (left) and 3D (right).